### v1

@@ Introduction
   Precision 0.78 0.70 0.90
   Recall 0.89 0.84 0.86
   F1-Score 0.83 0.76 0.88
@@
   Precision 0.52 0.08 0.52
   Recall 0.67 0.54 0.70
   F1-Score 0.59 0.14 0.60
@@
   Precision 0.80 0.69 0.81
   Recall 0.77 0.62 0.86
   F1-Score 0.78 0.65 0.84
@@
   Precision 0.96 0.95 0.94
   Recall 0.89 0.91 0.96
   F1-Score 0.92 0.93 0.95
Table 2: Comparison of BERT model performance trained on pure CONLL data, distilled data, and a combination of both
@@ Discussion
The hybrid training of small models with a mix of distilled data and human annotations presents higher adaptability for scenarios with limited resources or closed network environments. This adaptability is critical for applications where large language models cannot be directly accessed.
@@
The use of structured prompts outputting dictionaries, such as ’LOC’: [’Houston’], ’PER’: [’Orlando Miller’, ’Todd Stottlemyre’], ’ORG’: [’Houston Astros’, ’St. Louis Cardinals’, ’NL Central division’], enables the efficient representation of all entity types in a single request. This format can easily handle non-continuous entities and nested entities (e.g., "national basketball of America" as an made up organization containing "America" as a location), and it facilitates automated processing through conversion to dictionary objects. Additionally, the reasoning process uncovers certain annotation issues within the CONLL dataset, where GPT’s classifications sometimes appear more logical than the original labels.
@@ Conclusion
In conclusion, our research corroborates the significant potential of integrating Large Language Models (LLMs) such as GPT-4 into traditional NLP tasks, particularly Named Entity Recognition (NER). By employing the Chain of Thought (CoT) technique, we have successfully distilled knowledge from LLMs and utilized it to enhance the training of smaller models like BERT.
Our mixed-training strategy, which combines LLM-generated annotations with human annotations, has proven to be more effective than reliance on human efforts alone. This approach not only improves efficiency and performance but also reduces costs, aligning with our initial aim to leverage the strengths of LLMs while addressing their limitations in capacity and hallucination.
The application of our method is especially pertinent in resource-constrained and closed-network environments, showcasing the adaptability and practicality of smaller models when large-scale LLMs are not feasible.
Reflecting on our discussion, we recognize the challenges of LLM variability and the ongoing issues with hallucination. Future work will focus on advancing the sophistication of prompts, exploring parameterized prompting, and refining selection processes for LLM annotations to further improve the quality of NER models. Our continued efforts aim to optimize training strategies and expand the reach of this methodology across diverse NLP tasks.

### v3

@@ Introduction
(support: 2132)
recall 0.836 0.841 0.858 0.839 0.892 0.889 0.895 0.897 0.907 0.913
f1-score 0.687 0.688 0.725 0.719 0.817 0.818 0.834 0.827 0.844 0.847
precision 0.738 0.739 0.752 0.758 0.834 0.836 0.847 0.849 0.868 0.859
ORG
(support: 2669)
recall 0.644 0.644 0.701 0.685 0.802 0.802 0.822 0.806 0.822 0.836
f1-score 0.934 0.928 0.925 0.932 0.942 0.943 0.952 0.949 0.953 0.953
precision 0.941 0.927 0.930 0.943 0.934 0.939 0.951 0.947 0.952 0.953
PER
(support: 2768)
recall 0.927 0.929 0.921 0.922 0.951 0.947 0.953 0.951 0.955 0.954
f1-score 0.251 0.230 0.277 0.257 0.649 0.640 0.654 0.627 0.665 0.664
precision 0.167 0.150 0.191 0.169 0.613 0.603 0.611 0.588 0.614 0.618
@@
(support: 1029)
recall 0.504 0.500 0.508 0.548 0.690 0.684 0.706 0.673 0.727 0.718
Table 2: The comparison between different groups at phase two
simple mix ALL sigmoid(k=2) sigmoid(k=4) sigmoid(k=8) sigmoid(k=16) sigmoid(k=32) power(n=0.1) power(n=0.2) power(n=0.5) power(n=1) power(n=2) power(n=5) power(n=10) cosine
no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay no LR decay LR decay
f1-score 0.869 0.869 0.792 0.800 0.811 0.814 0.831 0.824 0.847 0.846 0.861 0.860 0.865 0.864 0.859 0.856 0.861 0.861 0.846 0.847 0.848 0.849 0.840 0.852 0.855 0.845 0.849 0.849 0.853 0.854
precision 0.864 0.859 0.747 0.757 0.773 0.782 0.811 0.796 0.825 0.823 0.846 0.845 0.851 0.853 0.850 0.839 0.846 0.846 0.826 0.826 0.826 0.831 0.819 0.834 0.841 0.826 0.839 0.842 0.834 0.836
micro avg
(support: 8598)
recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
f1-score 0.840 0.839 0.755 0.762 0.774 0.773 0.792 0.786 0.810 0.809 0.828 0.827 0.835 0.833 0.827 0.823 0.828 0.826 0.811 0.811 0.813 0.812 0.805 0.817 0.822 0.808 0.816 0.812 0.819 0.818
precision 0.830 0.825 0.727 0.734 0.745 0.747 0.776 0.761 0.786 0.786 0.810 0.809 0.818 0.818 0.814 0.804 0.809 0.809 0.790 0.788 0.788 0.793 0.782 0.797 0.806 0.788 0.805 0.802 0.798 0.799
macro avg
(support: 8598)
recall 0.853 0.855 0.813 0.818 0.826 0.815 0.821 0.826 0.840 0.840 0.851 0.850 0.855 0.850 0.842 0.847 0.851 0.847 0.842 0.841 0.846 0.839 0.836 0.844 0.843 0.835 0.830 0.825 0.848 0.844
f1-score 0.871 0.871 0.814 0.821 0.827 0.827 0.839 0.835 0.851 0.852 0.864 0.864 0.867 0.866 0.862 0.860 0.864 0.864 0.852 0.853 0.854 0.855 0.845 0.857 0.858 0.849 0.850 0.851 0.858 0.858
precision 0.869 0.864 0.801 0.808 0.813 0.813 0.833 0.824 0.836 0.838 0.854 0.855 0.858 0.858 0.855 0.848 0.855 0.854 0.843 0.841 0.841 0.846 0.833 0.845 0.848 0.838 0.843 0.847 0.847 0.848
weighted avg
(support: 8598)
recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
f1-score 0.897 0.892 0.855 0.858 0.864 0.864 0.868 0.869 0.881 0.884 0.895 0.890 0.890 0.893 0.890 0.889 0.888 0.895 0.883 0.884 0.883 0.884 0.871 0.883 0.888 0.874 0.882 0.876 0.886 0.891
precision 0.888 0.872 0.828 0.827 0.833 0.840 0.864 0.842 0.854 0.869 0.878 0.877 0.871 0.875 0.880 0.879 0.871 0.884 0.859 0.870 0.859 0.867 0.843 0.868 0.883 0.853 0.868 0.862 0.871 0.878
LOC
(support: 2132)
recall 0.907 0.913 0.884 0.893 0.898 0.889 0.874 0.897 0.909 0.900 0.913 0.903 0.910 0.912 0.900 0.900 0.905 0.906 0.909 0.898 0.908 0.902 0.903 0.900 0.894 0.896 0.898 0.891 0.903 0.904
f1-score 0.844 0.847 0.797 0.804 0.810 0.808 0.823 0.815 0.831 0.832 0.842 0.846 0.841 0.841 0.839 0.836 0.843 0.849 0.828 0.837 0.838 0.837 0.823 0.837 0.835 0.832 0.822 0.830 0.837 0.837
precision 0.868 0.859 0.822 0.832 0.834 0.813 0.835 0.840 0.834 0.824 0.854 0.854 0.845 0.852 0.848 0.832 0.851 0.846 0.848 0.846 0.853 0.850 0.831 0.836 0.846 0.843 0.833 0.843 0.843 0.844
@@
(support: 2669)
recall 0.822 0.836 0.776 0.778 0.789 0.803 0.813 0.792 0.829 0.842 0.830 0.839 0.839 0.831 0.831 0.841 0.836 0.853 0.810 0.828 0.825 0.825 0.816 0.838 0.825 0.822 0.813 0.817 0.832 0.830
f1-score 0.953 0.953 0.946 0.950 0.947 0.948 0.949 0.948 0.950 0.950 0.950 0.952 0.954 0.949 0.947 0.950 0.956 0.950 0.952 0.949 0.951 0.953 0.945 0.951 0.943 0.946 0.939 0.948 0.952 0.952
precision 0.952 0.953 0.942 0.947 0.941 0.950 0.948 0.945 0.945 0.956 0.943 0.949 0.957 0.946 0.942 0.948 0.956 0.950 0.953 0.942 0.944 0.954 0.946 0.951 0.929 0.940 0.928 0.949 0.950 0.949
@@
(support: 2768)
recall 0.955 0.954 0.951 0.954 0.953 0.946 0.951 0.950 0.956 0.945 0.958 0.954 0.952 0.952 0.952 0.952 0.955 0.950 0.951 0.955 0.957 0.951 0.944 0.951 0.958 0.953 0.952 0.948 0.954 0.955
f1-score 0.665 0.664 0.421 0.437 0.475 0.474 0.527 0.513 0.577 0.570 0.626 0.620 0.654 0.650 0.631 0.618 0.625 0.611 0.581 0.573 0.578 0.575 0.582 0.597 0.623 0.581 0.620 0.595 0.601 0.593
precision 0.614 0.618 0.315 0.331 0.372 0.383 0.457 0.417 0.508 0.496 0.564 0.555 0.599 0.602 0.586 0.557 0.559 0.555 0.499 0.494 0.497 0.499 0.509 0.531 0.567 0.517 0.590 0.555 0.527 0.523
MISC
(support: 1029)
recall 0.727 0.718 0.640 0.648 0.666 0.624 0.645 0.667 0.668 0.674 0.703 0.705 0.721 0.706 0.684 0.697 0.709 0.679 0.698 0.683 0.693 0.679 0.680 0.686 0.694 0.668 0.659 0.643 0.702 0.686
Table 3: Comparison of performance between strategies
@@ Discussion
The hybrid training of small models with a mix of distilled data and human annotations presents higher adaptability for scenarios with limited resources or closed network environments. This adaptability is critical for applications where large language models cannot be directly accessed. Our experiments also reveal that incorporating distilled data from external sources related to the domain of interest (e.g., BBC news data, in addition to CONLL2003) can lead to performance improvements. This supports the notion that leveraging similar data types for distillation is beneficial, as reflected in our NER task results.
@@
The use of structured prompts outputting dictionaries, such as ’LOC’: [’Houston’], ’PER’: [’Orlando Miller’, ’Todd Stottlemyre’], ’ORG’: [’Houston Astros’, ’St. Louis Cardinals’, ’NL Central division’], enables the efficient representation of all entity types in a single request. This format can easily handle non-continuous entities and nested entities, and it facilitates automated processing through conversion to dictionary objects. Additionally, the reasoning process uncovers certain annotation issues within the CONLL dataset, where GPT’s classifications sometimes appear more logical than the original labels.
@@
While the ’Simple Mix’ strategy yielded the most effective results in our experiments, the underlying mechanisms contributing to its success are not fully understood. Future research will delve deeper into this phenomenon to uncover the reasons behind its efficacy and explore whether this pattern holds across different LLMs, datasets, and NLP tasks.
@@ Conclusion
In conclusion, this research substantiates the transformative potential of integrating Large Language Models (LLMs) like GPT-4 into traditional NLP tasks, particularly Named Entity Recognition (NER). By harnessing the Chain of Thought (CoT) prompting, we have effectively distilled LLM knowledge to enhance the training of more compact models such as BERT.
Our investigation has solidified the efficacy of a mixed-training strategy that synergizes LLM-generated annotations with human annotations. This combined approach surpasses the performance of models trained solely on human-annotated data, leading to improvements in efficiency, cost-effectiveness, and model accuracy—goals we set out to achieve by leveraging LLM strengths and mitigating their limitations.
Our method has demonstrated particular value in resource-limited and closed-network settings, providing a versatile solution where the deployment of large-scale LLMs is not viable. The adaptability and practicality of smaller models are thus accentuated, offering a viable pathway to harnessing the power of LLMs in a broader range of contexts.
Reflecting on the challenges encountered, such as LLM variability and hallucination tendencies, we see clear directions for future work. Our research will extend to enhance prompt sophistication, investigate parameterized prompting, and refine the selection of LLM annotations to further elevate NER model quality. We will also delve deeper into optimizing training strategies, as the ’Simple Mix’ approach showed the most promise in our studies. Expanding upon these findings, our future endeavors will aim to refine data blending techniques and explore their applicability across a spectrum of NLP tasks, with the goal of continuously pushing the boundaries of what’s achievable in the field of natural language processing.